OpenStudy (anonymous):

Rearrange the formula by copying and completing each line. 2y-6x-12=0 2y-12=______ 2y=______ y=______

OpenStudy (texaschic101):

2y - 6x - 12 = 0 2y - 12 = 6x 2y = 6x + 12 y = (6x + 12)/2 OR y = 3x + 6

OpenStudy (texaschic101):

oh...lol....I kinda like your username :) Its funny

OpenStudy (anonymous):

Haha thanks

OpenStudy (anonymous):

i understand now